### v1

# GP-MoLFormer: A Foundation Model For Molecular Generation

Abstract
Transformer-based models trained on large and general purpose datasets consisting of molecular strings have recently emerged as a powerful tool for successfully modeling various structure-property relations. Inspired by this success, we extend the paradigm of training chemical language transformers on large-scale chemical datasets to generative tasks in this work. Specifically, we propose GP-MoLFormer, an autoregressive molecular string generator that is trained on more than 1.1B chemical SMILES. GP-MoLFormer uses a 46.8M parameter transformer decoder model with linear attention and rotary positional encodings as the base architecture. We explore the utility of GP-MoLFormer in generating novel, valid, and unique SMILES. Impressively, we find GP-MoLFormer is able to generate a significant fraction of novel, valid, and unique SMILES even when the number of generated molecules is in the 10 billion range and the reference set is over a billion. We also find strong memorization of training data in GP-MoLFormer generations, which has so far remained unexplored for chemical language models. Our analyses reveal that training data memorization and novelty in generations are impacted by the quality of the training data; duplication bias in training data can enhance memorization at the cost of lowering novelty. We evaluate GP-MoLFormer’s utility and compare it with that of existing baselines on three different tasks: de novo generation, scaffold-constrained molecular decoration, and unconstrained property-guided optimization. While the first two are handled with no additional training, we propose a parameter-efficient fine-tuning method for the last task, which uses property-ordered molecular pairs as input. We call this new approach pair-tuning. Our results show GP-MoLFormer performs better or comparable with baselines across all three tasks, demonstrating its general utility for a variety of molecular generation tasks.
Main
Identifying molecules with desirable properties from the vast landscape of possibilities is daunting. It has been postulated that only $10^{8}$ out of $10^{23}$$10^{60}$ possible small drug-like molecules are synthesizable [1]. As the search space is enormous and high-throughput screening of molecules is costly and time-consuming, this requires a thorough understanding of the chemical data manifold. Recently, similar to what has been experienced in computer vision and natural language processing, deep generative models have made great strides in modeling molecular distributions and sampling new molecules from them in de novo or targeted manners. Among those efforts, a significant fraction use string-based representations of molecules as the input; thus, techniques explored in language modeling, such as causal and masked language modeling, are becoming widely used in building molecular deep neural models.
Interestingly, much of the recent performance gains for natural language models have come from scale — in terms of the number of parameters and the number of training samples [2, 3, 4]. It is also clear that larger language models that can memorize training data show improved generalization [5, 6, 7]. However, perhaps surprisingly, relatively less work has been done in understanding the impact of scale and the impact of training data memorization on the performance of generative models of molecules. Specifically, it remains under-explored to what extent a causal large language model of molecules, trained on large-scale (>100M) training data, memorizes its training data and demonstrates such memorization in its generations.
In the context of memorization of language models, it is now known that data that is seen during training many times is memorized more, and de-duplication of training data plays a big role in preventing such memorization  [5, 6, 8]. In chemical language modeling tasks, molecules in training data originate from publicly available databases such as ZINC [9] and PubChem [1]. It is known that certain molecules, as well as certain molecular features, are over-represented in those databases [10]. How such bias is perpetuated by generative chemical language models remains relatively unknown.
To bridge these gaps, in this work, we present a family of generative pre-trained molecular foundation models for the unconstrained and targeted generation of novel molecules, which are based on the recently published Molecular Language transFormer (MoLFormer) architecture [11]. We refer to these Generative Pre-trained models as GP-MoLFormer. The base transformer architecture of our decoder-only GP-MoLFormer, consisting of $\approx 47$M parameters, uses an efficient linear attention mechanism, together with rotary positional encodings, analogously to MoLFormer [11]. The model is then trained with a causal language modeling objective on a large corpus of 0.65–1.1 billion canonicalized SMILES strings of small molecules from publicly available chemical databases. We evaluate memorization in GP-MoLFormer, in the light of the bias present in training data, and how such memorization affects the model’s generation performance in the de novo molecule generation task. To our knowledge, this is the first report on effect of training data memorization in a generative pre-trained chemcial language model. Experiments demonstrate that GP-MoLFormer is able to generate novel, unique, diverse, and valid molecules even when the generation pool reaches a size of 10B while showing consistent memorization of training data. We further explore how representational bias encoded in the public chemical databases is perpetuated by a generative chemical language model and is reflected in its generations.
We further extend GP-MoLFormer to two targeted molecular design tasks: scaffold-constrained molecular decoration and unconstrained property-guided optimization. For the first, we exploit GP-MoLFormer’s causal language modeling ability and establish GP-MoLFormer’s ability to handle the task without undergoing any task-specific tuning. For the second, we provide a prompt-tuning or soft prompt-learning algorithm that learns from partial orderings of molecules. We name this method pair-tuning. Results show that pair-tuning on GP-MoLFormer provides on par or better performance in three different property optimization tasks, namely (i) drug-likeness optimization, (ii) penalized logP optimization, and (iii) optimization of dopamine type 2 receptor binding activity.
Results and Discussion
GP-MoLFormer Framework
We construct the GP-MoLFormer architecture (Figure 1A) using a 12-layer linear attention transformer with 12 attention heads and a hidden state size of 768 and generalized random Fourier feature maps [12] for the linear attention. The input representation considered is chemical SMILES. We choose a relative positional embedding, namely rotary positional embedding [13], to model the token dependency within SMILES better using rotary attention. During training, GP-MoLFormer uses a causal modeling objective of predicting the next token given the context history of prior tokens in the input SMILES strings. Given the size of the model and the efficient linear attention, GP-MoLFormer takes only around 3 milliseconds for a single forward pass during generation. After pre-training, we quantitatively assess the performance of GP-MoLFormer on the generation of de novo molecules before applying a novel prompt-tuning algorithm for molecular optimization (Figure 1B).
Figure 1: GP-MoLFormer— a generative pre-trained molecular foundation model. (A) Unconditional generation using GP-MoLFormer. SMILES representations are generated autoregressively and randomly along the learned manifold (purple area). (B) During pair-tuning, a prompt vector is learned, which translates a given molecular representation (light blue dots) to an optimized region of the manifold (red diamonds).
De novo generation of molecules
The task under consideration is to generate random molecules by sampling from the generative model. As downstream optimization can rely on these randomly generated molecules as starting points, the generated distribution must contain novel, diverse, and unique molecules.
Effect of training dataset and generation pool
The generative ability of a chemical language model can be affected by the scale of the training data. Further, bias in training data can also contribute to the quality of generation. To disentangle these factors, we report GP-MoLFormer’s performance on two independently varying dimensions: First, we vary the size and quality of pre-training data. For this dimension, we consider training the GP-MoLFormer on (i) 1.1B SMILES extracted from ZINC and PubChem, as described in Ross et al. (2022)[11], and (ii) a de-duplicated 650M subset of (i). GP-MoLFormer trained on (i) is referred hereafter GP-MoLFormer-Raw, while GP-MoLFormer trained on (ii) is termed as GP-MoLFormer-Clean. Secondly, we analyze the effect of varying the number of generated molecules from 30k to 10B molecules. It should be noted that existing molecular generation benchmarks consider reporting performance metrics for a relatively small 10–30k generations, as those models are designed to target a specific distribution of molecules, e.g., synthetic molecules with biological activity or natural products, and are trained on 1–100M samples[14]. In contrast, GP-MoLFormer is trained on 650M–1.1B SMILES and the dataset captures the relative abundance of molecules, as well as the presence of the same molecule in different context, as found in chemical databases.
We report in Table 1 the percentage of novel (unseen in training), valid (syntactically correct), and unique (not previously generated) molecules for both GP-MoLFormer-Raw and GP-MoLFormer-Clean, for generation size of 30k to 10M. The results suggest that the fraction of novel generations stays at a consistent $\approx$ 32% for GP-MoLFormer-Raw when the number of total generated molecules is within this range. Novelty in GP-MoLFormer-Clean is $\approx$7–8higher compared to that of GP-MoLFormer-Raw for all generation pool sizes and rises with the increasing pool size. GP-MoLFormer, irrespective of training data, outputs chemically valid SMILES almost all the time. While the percentage of unique molecules drops slightly with increasing generation pool size, it still is over 99% for 10M generations. This result confirms that (i) GP-MoLFormer memorizes training data, as seen from the high number of exact matches ($1-$novelty, which can be up to 60%) with training molecules; and (ii) memorization is less when the training data is de-duplicated, enabling more novel generation. Nevertheless, in all cases studied here, GP-MoLFormer is capable of generating novel, diverse, and valid molecules.
We continue exploring the quality of generations while increasing the number of samples generated to billions for GP-MoLFormer-Raw. Results are shown in Table 2. At or beyond 1B generations, the fraction of novel and unique generations drops but still remains significant. Even for 10B generations, GP-MoLFormer-Raw is able to generate a significant 16.7% novel molecules. And all throughout, 99% of generations are chemically valid.
Training Size = 650M Training Size = 1.1B
Generation Size Novel Unique Valid Novel Unique Valid
30k 0.390 0.997 1.000 0.323 0.997 0.997
100k 0.393 0.996 0.999 0.326 0.998 0.998
1M 0.395 0.996 0.999 0.323 0.996 0.997
10M 0.400 0.991 0.994 0.322 0.989 0.997
Table 1: Novelty, validity, and uniqueness of different number of generations for GP-MoLFormer trained with 650M (Clean) and 1.1B (Raw) training set.
Size Novel Unique Valid
30k 0.323 0.997 0.997
100k 0.326 0.998 0.998
1M 0.323 0.996 0.997
10M 0.322 0.956 0.997
100M 0.327 0.989 0.997
1B 0.278 0.611 0.997
10B 0.167 0.223 0.997
Table 2: Metrics on GP-MoLFormer-Raw generations of different scale.
Effect of sampling temperature
We also report the generative performance of GP-MoLFormer-Raw at several softmax temperature values, as we use multinomial sampling during decoding to encourage novel generation. In fact, recent work has reported the utility of multinomial sampling over beam search for generating molecules that are simultaneously novel and useful (in terms of synthesis and biological testing) [15]. Results are shown in Figure 2, which shows that at around $T=1.2$, GP-MoLFormer-Raw performs the best in terms of balancing high novelty (47%), uniqueness (99%), and validity (98%). Further, when we compare novelty of the same generated sets with respect to a random subset of 130M molecules from training set, we see novelty is significantly higher (around 90% at $T=1.2$), showing the effect of training data scale on generation novelty — there are simply more opportunities to match with a larger reference set. All results elsewhere in this paper are generated using $T=1.0$.
Figure 2: Effect of temperature during multinomial sampling of 100k generated molecules on the percentage of novel molecules (blue) with respect to full 1.1B training set, novelty with respect to a random 130M subset of the training (orange), validity (green), and uniqueness (red) of generations from GP-MoLFormer-RAW.
Discussions on the effect of training data bias on generations
Within the Raw training set, a notable 45% of SMILES strings were found duplicates. Some of this is due to molecules that appear multiple times as different isomers, which are pre-processed into the same SMILES representation in our canonicalization pipeline. There is also some overlap between the two databases — ZINC and PubChem. This popularity of certain molecules may or may not be related to true factors, such as the molecule being useful in a multitude of applications because of its synthetic ease or lower cost, or a combination of all of those. It is reported that such skewed distribution can also originate from anthropogenic attention bias [10], where some molecules are studied extensively because scientists themselves or their peers “like” them. Nevertheless, such overrepresenation of certain molecules, either originating from human cognitive biases, heuristics, social influences, from rationally made choices, from multiple isomeric instances, or a combination of all of the above, can lead to bias in databases.
Data de-duplication is the first step towards removing such bias, which reduces the concentration of the high density regions of the data manifold. In this case, de-duplication removes isomeric population information as well as repeated molecules across databases. The de-duplicated data is closer to a data manifold that has more homogenized density all over. Training on such a data manifold results in higher novelty in generations, as found for GP-MoLFormer-Clean when compared to GP-MoLFormer-Raw, as shown in Table 1.
While existing molecular generative models trained on a much smaller and much focused datasets have demonstrated near-perfect (100%) novelty in generations, they are for most part is not suitable for studying the trade-off between training data memorization and generation novelty. Investigating such phenomenon requires studying a generative chemical (language) model that has been trained on a broader-purpose and much larger dataset at scale. Our experiments attempt to address this under-explored aspect in this study. As shown in Table 1, novelty in GP-MoLFormer generations are lower compared to $\approx 100\%$ reported by baselines[14], but still sufficiently high for practical use. When compared with recent baselines, GP-MoLFormer generations are more dissimilar to test molecules (Table 3), though GP-MoLFormer’s test set is more diverse. And, finally, the low novelty in GP-MoLFormer’s generations is reflective of modeling its training set that represents the relative usage of molecules in real-world.
Similarly, the present study highlights the importance of studying generated sets of different sizes to obtain a comprehensive view of the quality of generations, particularly when the generative model is trained on data at scale. As GP-MoLFormer-Clean aims to capture a training data manifold of more uniform density, which is enabled by de-duplicating the training SMILES, we see a 1% rise in novelty as we increase the number of generated samples from 30k to 10M. A similar observation has been reported in image generation [16] and language generation [17, 8]. To summarize, novelty in generations is influenced by the support provided by both the training distribution and the generated distribution, and should therefore be assessed relative to the sizes and diversity of those two sets.
These results in Table 1 complement and support earlier efforts focusing on studying scaling behaviors of chemical language models. One such noteworthy effort along this line is Frey et al. (2023)[18], where neural-scaling behaviour in large chemical models was investigated by studying models with over 1B parameters. However, the models tested in that work were only pre-trained on datasets of size up to $\approx$ 10M data points, which is very small compared to the size of the chemical universe. In Ross et al. (2022)[11], the scaling behavior of MoLFormer, which is a transformer-based molecular encoder built using a masked language modeling objective, was studied. That work clearly established the scaling behavior underlying adaptation of a pre-trained model across downstream tasks, in which the number of model parameters was up to 47M while the number of training points considered was >1B. It was shown that a MoLFormer trained on 100M SMILES consistently underperformed across a wide variety of property prediction tasks, including quantum mechanical and physiological, when compared to the model trained on >1.1B SMILES, indicating predictive ability may benefit from such bias in training data. In contrast, the results in Table 1 show that a generative chemical language model trained on cleaner de-duplicated training data produces more novel generations.
Comparison of GP-MoLFormer with baselines
Next, we compared GP-MoLFormer-Clean with different baseline models, such as the character-level recurrent neural network (CharRNN) [14], SMILES variational autoencoder (VAE) [14], junction tree VAE (JT-VAE) [19], latent inceptionism on molecules (LIMO) [20], and MolGen-7b [21]. We consider baselines that are not fine-tuned to particular molecular domains. Except MolGen-7b, all baselines were trained and tested on datasets from MOSES[14], whose origin is the ZINC Clean Leads dataset[9]. The size of that training set is 1.6M. MolGen-7b was trained on 100M filtered molecules from ZINC-15[9] as detailed in Irwin et al. (2022)[22]. LIMO and MolGen-7b are trained using an alternative molecular string representation, SELFIES [23], that guarantees 100% validity of generated molecules. We, in contrast, train GP-MoLFormer on SMILES as recent work shows that training a generative language model on SELFIES may hurt model’s exploratory ability [24]. All baseline performances are reported on their corresponding test set consisting of 175k molecules (if the original test set was larger, this is a randomly selected subset). While GP-MoLFormer-Clean’s performance is estimated on a held-out test set that is of similar size, we found this test set to be more diverse in terms of number of unique scaffolds present within the set (126k compared to 124k in the ZINC-15 subset and 77k in the MOSES set) and by comparing different property distributions with that of the other baselines (see Figure 3).
Standard metrics for evaluating model-generated molecules are reported in Table 4 for a generation set of 30k molecules. When compared to baselines, GP-MoLFormer-Clean is equally performant in generating unique, valid, and novel molecules that share high cosine similarity with the corresponding reference molecules at the fragment (Frag) level, consistent with low Fréchet ChemNet Distance (FCD)[25]. At the same time, GP-MoLFormer-Clean generates molecules with high internal diversity (IntDiv), i.e., average pairwise dissimilarity. The scaffold cosine similarity (Scaf) and similarity to the nearest neighbor in the test set (SNN) of GP-MoLFormer-Clean is comparable to that of baselines for 30k generations. All metrics are computed using MOSES[14].
In order to quantify the amount of novelty in generations, we also calculated the pairwise similarity between novel and unique generations and molecules from the corresponding 175k sample test set. We then report both the average similarity per generated molecule and the maximum similarity per generated molecule over the test set. These results are presented in Table 3. GP-MoLFormer-Clean is slightly lower than MolGen-7b in both average mean and average maximum similarity, indicating generations that are more dissimilar with respect to its test set. LIMO is much lower than both of these, though, as we see in Table 4, the outputs of this model do not match its test set well, so this is to be expected. Taken together, this means GP-MoLFormer-Clean is able to generate molecules that are about as novel as other models, despite the seemingly low novelty scores in Table 1 (see also, the discussion of Figure 2).
Model mean max
LIMO 0.0905 0.2474
MolGen-7b 0.1385 0.5138
GP-MoLFormer-Clean 0.1354 0.4533
Table 3: Average Tanimoto similarity between 30k randomly generated molecules and all respective test molecules (mean) or most similar test molecule (max).
We next investigate how these metrics change with varying number of generated and test molecules. Table 5 shows that, with increasing the generated pool size, scaffold similarity with respect to the test molecules becomes $>0.9$ while SNN reaches $>0.5$ when compared against 175k held-out test samples. When a larger test set of 1M molecules is used, further increases in both scaffold similarity and SNN are observed. These results suggest that GP-MoLFormer is able to model the training data faithfully. Further, our work implies that, with increasing diversity of the training data, the typical metrics used in assessing molecular generative models, such as various similarity measures with respect to a test set, should be analyzed with generation and test sets that are larger in size compared to what is typically used in the field.
Figure 3: Property distributions of different test datasets — MOSES, ZINC-15 (MolGen-7b), and GP-MoLFormer-Clean (ours). Clockwise from top left: octanol-water partition coefficient, drug-likeness, synthetic accessibility, molecular weight. Our test distributions are consistently wider (more diverse) than the other baselines.
[…]
CharRNN [14] 0.9998 0.9242 0.6015 0.8562 0.0732
VAE [14] 0.9984 0.9386 0.6257 0.8558 0.0990
JT-VAE [19] 0.9965 0.8964 0.5477 0.8551 0.3954
LIMO [20] 0.6989 0.0079 0.2464 0.9039 26.78
MolGen-7b [21] 0.9999 0.6538 0.5138 0.8617 0.0435
GP-MoLFormer-Clean 0.9998 0.7383 0.5045 0.8655 0.0591
Table 4: Comparison of 30k generations with a held-out test set of size 175k. Baseline performances for CharRNN, VAE, and JT-VAE are taken from MOSES[14]. LIMO is reproduced using their random generation model trained on the MOSES data. These models all use the default MOSES test split for reference. The MolGen-7b baseline uses the pre-trained model from Hugging Facehttps://huggingface.co/zjunlp/MolGen-7b with multinomial sampling ($T=1.0$) and is tested on a random 175k subset of the original test data. GP-MoLFormer-Clean is tested with respect to a held-out 175k set from its training data.
Test size Gen. size Frag Scaf SNN IntDiv FCD
175k 30k 0.9998 0.7383 0.5045 0.8655 0.0591
100k 0.9998 0.8653 0.5045 0.8657 0.0279
1M 0.9998 0.9375 0.5040 0.8658 0.0178
1M 30k 0.9998 0.7702 0.5738 0.8655 0.0646
100k 0.9998 0.9026 0.5740 0.8657 0.0331
1M 0.9998 0.9786 0.5739 0.8658 0.0227
Table 5: Investigation on the interplay between generation set size and test set size on modeling the molecular distribution.
Scaffold-constrained decoration
We further subject GP-MoLFormer to the task of scaffold-constrained generation. For this experiment, we first take the five unique scaffolds from the dopamine receptor $\mathrm{D}_{2}$ (DRD2) active binder dataset validation split[26]. These scaffolds contain between two to four attachment points. We perform a pre-processing step for each unique scaffold, generating every possible randomized SMILES representation of that scaffold. Then we sort the resulting candidates according to the distance of the “*” characters, also known as the attachment point, to the end of the string. For multiple *s, we sum the distances. As an example, C1(=O)N(CCN1*)* would score $2(2+0)$ while C1N(*)C(=O)N(*)C1 would score $15(12+3)$. If multiple representations are equivalently optimal, we save all of them. During the generation step, we provide the candidates produced in this pre-processing step as input to GP-MoLFormer.
Next, the task is to generate multiple possible candidates for the first attachment point given an input scaffold. First, we collect all valid candidates from that generation. Then, we again generate multiple possible candidates for the recently extended scaffolds. This process is repeated until all the attachment points are decorated then we collect all valid molecules generated.
We compare the performance of GP-MoLFormer in terms of generating DRD2 active molecules that will pass the DRD2 binding classifier ($p>0.5$). For baselines of comparison, we consider our own random generations from GP-MoLFormer-Raw, as well as an earlier scaffold-conditioned generation model [26] that was specifically trained for scaffold decoration tasks and was then used to decorate the same scaffolds under investigation here with fragments from ChEMBL. In contrast to this baseline model, GP-MoLFormer-Raw has not seen scaffold-constrained generation during pre-training, nor is it specifically fine-tuned for this purpose. Table 6 shows that GP-MoLFormer-Raw generates more DRD2 active hits compared to a random baseline of de novo generation, as well as a generative model trained on this specific task.
predicted active hits (%)
Scaffold decorator[26] 3.64
de novo GP-MoLFormer-Raw 0.83
scaffold-conditioned GP-MoLFormer-Raw 4.58
Table 6: Scaffold-constrained generation. Baseline performance is taken from Arús-Pous et al. (2020)[26].
Unconstrained Single Property Optimization
Given GP-MoLFormer demonstrates desirable performance in both novel molecule generation and scaffold-constrained molecular decoration, it makes sense to extend GP-MoLFormer to downstream task settings, where the goal is to generate molecules with a desired property. In light of current LLM adaptation efforts, one obvious path is model tuning (or “fine-tuning”), where all model parameters are tuned during adaptation. This approach often is highly data-hungry. As an alternative, prompt-tuning or “soft prompt” learning has been proposed, which includes an additional $n$ tuneable tokens for each downstream task, which is prepended to the input text [27]. This soft prompt is then trained end-to-end on a labeled dataset, whereas the pre-trained LLM remains frozen. This method has been demonstrated to close the gap in model tuning, even when combined with a smaller LLM [27] and is lower cost compared to full fine-tuning.
We exploit prompt-tuning to introduce a novel means for enabling GP-MoLFormer to tackle property-specific molecular optimization tasks, where the goal is to generate molecules with a specific property value above a pre-defined threshold. Below, we describe the pair-tuning framework and then show that pair-tuning performs well on a set of three tasks.
Pair-tuning framework
Following a “text-to-text” approach, we formalize the task of generating a (property-)optimal molecule as follows: Given a molecule $a$, translate it to another molecule $b$ with a more optimal property value where $a,b$ come from domain $\Omega$. This conditional generation task is $P_{\theta}(b|a)$, where $\theta$ is the parametrization of the generative language model. This task is handled via learning soft prompts, i.e., prompt-tuning, which is a parameter-efficient task adaptation method for a frozen language model [27]. Specifically, we add a small number of task-specific parameters $\phi_{T}$, such that the conditional task becomes $P_{\theta}(b|\phi_{T},a)$ and is trained through maximizing the probability likelihood of $b$. Only $\phi_{T}$ is updated during gradient backpropagation. This procedure is explained in Algorithm 1.
In this formulation, we do not need absolute property values of the molecules, rather only ordered pairs of molecules are needed. This is to mimic the scenario of many drug and material development tasks, in which two molecules are compared with each other to guide molecular optimization and prioritization, especially for tasks with limited available data. For example, Matched Molecular Pair (MMP) analysis allows the rapid estimation of property differences [28, 29]. However, MMP analysis is limited to comparing close molecular derivatives and common molecular derivations, and it can fail to model important chemical contexts. The present formulation of optimizing molecules is free from such constraints and only aims to learn task-specific soft prompts to generate more optimal molecules given a seed molecule.
Algorithm 1 Pair-tuning training
1:Pretrained GP-MoLFormer; Number of fine-tuning epochs, $m$; number of enhancement tokens, $n$; Pair-Tuning dataset $D=\{(a,b)|b>a\}$ for $a,b\in\Omega$ where $>$ is an order relation defined for a certain property $T$
2:Append $n+1$ new tokens $\phi_{T}=\langle enh_{1}\rangle\langle enh_{2}\rangle\dots\langle enh_{n}\rangle,\langle sep\rangle$ to GP-MoLFormer-Raw vocabulary
3:for $m$ epochs do
4:     Prepare training prompts as $\langle enh_{1}\rangle\langle enh_{2}\rangle\dots\langle enh_{n}\rangle\langle bos\rangle\langle a_{1}\rangle\langle a_{2}\rangle\dots\langle a_{i}\rangle\langle sep\rangle$ where $a=\langle a_{1}\rangle\langle a_{2}\rangle\dots\langle a_{i}\rangle$, i.e., tokenized training molecules
5:     Compute the cross-entropy (CE) loss conditioned on enhancement tokens $\phi_{T}$ and molecule $a$ with the auto-regressive CE loss with target the molecule $b=\langle b_{1}\rangle\langle b_{2}\rangle\dots\langle b_{n}\rangle\langle eos\rangle$
6:     Compute the gradient of the auto-regressive CE loss with respect to enhancement tokens $\phi_{T}$
7:     Update enhancement tokens via gradient descent optimizer
8:end for
Pairtuning tasks
We evaluate pairtuning using GP-MoLFormer-Raw on three property optimization benchmarks, namely drug-likeness (QED) maximization, penalized logP maximization, and activity maximization for DRD2. The first two properties, QED and penalized logP, are important considerations in drug discovery, and these task shows the ability of a model to optimize salient aspects of a molecule, even if maximization of these properties by themselves is of low utility[30]. The goal of the third task is to increase the binding affinity of a compound to a ligand, in this case, the dopamine $\mathrm{D}_{2}$ receptor.
Penalized $\log\mathrm{P}$ QED
1 st 2nd 3rd Validity $1\mathrm{st}$ 2nd 3rd Validity
JT-VAE 5.30 4.93 4.49 100% 0.925 0.911 0.910 100%
MARS 45.0 44.3 43.8 100% 0.948 0.948 0.948 100%
GRAPHDF 13.7 13.2 13.2 100% 0.948 0.948 0.948 100%
LIMO on $z$ 6.52 6.38 5.59 100% 0.910 0.909 0.892 100%
LIMO 10.5 9.69 9.60 100% 0.947 0.946 0.945 100%
$\mathrm{GCPN}$ 7.98 7.85 7.80 100% 0.948 0.947 0.946 100%
MolDQN-bootstrap 11.84 11.84 11.82 100% 0.948 0.944 0.943 $100\%$
Pair-tuning ($k=125$) 13.18 (7.12) 12.24 (6.61) 11.51 (6.40) 94.7% 0.948 0.947 0.947 94.7%
Pair-tuning ($k=1000$) 19.59 (9.35) 15.51(8.93) 15.27 (8.64) 94.5% 0.948 0.948 0.948 94.5%
Table 7: Performance on unconstrained penalized logP and QED optimization. Baseline performances are taken from Zhou et al. (2019)[30] and Eckmann et al. (2022)[20] and are reported on 100k generations as per LIMO[20]. For GP-MoLFormer, we set $k$ the number of targeted generation attempts per molecule to $125$: given a test set of size 800 this results in 100k total generations.
Penalized logP optimization
Table 7 shows results of pair-tuning on GP-MoLFormer-Raw, as well as of the baselines, in terms of the generated molecules with high penalized logP. Penalized logP is calculated as $\log\mathrm{P}-\mathrm{SA}-{\max(\max_{\mathrm{rings}}(\mathrm{size})-6,0)}$, i.e., logP penalized by SA and maximum ring size, if larger than 6. We report pair-tuning performances as a function of two different $k$, where $k$ is the number of targeted generation attempts per molecule. For $k=125$, using a test set containing 800 molecules gives a total number of generated molecules of 100k, which is the same used for the baselines. The baselines under consideration are JT-VAE [19], GCPN [31], MolDQN [30], MARS [32], GraphDF [33], and LIMO [20]. Penalized logP can be artificially inflated simply by generating molecules with increased length, specifically by adding alkyl carbons [32, 20]. Many works, e.g., GCPN, MolDQN, and LIMO, avoid this by reporting top property scores given length constraints, e.g., limiting the length up to the maximum molecule size of the ZINC250k dataset. MARS, on the other hand, does not consider such a length constraint. We also report the top 3 scores for pair-tuning with a length constraint (length < 38), added post generation, in Table 7 as the value within parentheses.
Compared to the strongest length-constrained baselines, pair-tuning generates molecules with comparably high values. When the length constraint is not considered, pair-tuning still generates molecules with higher but reasonable penalized logP values. Note that pair-tuning does not require feedback or evaluation on generations from an additional reward model or a property predictor, nor is the generative model updated during the tuning. We also report top 3 scores for 1M generations ($k=1000$), which requires less than an hour to generate. Although all the baselines produce molecules with 100% validity due to their methods utilizing SELFIES or graphs, our method’s validity is still very high (around 95%) and overall this is negligible compared to the ease of generating additional molecules. Altogether, Table 7 shows that the proposed method can generate molecules with even higher penalized logP values, both with and without a length constraint.
QED optimization
As with penalized logP, we show results for QED optimization in Table 7 compared with the same baselines. Again, pair-tuning performances are reported for two different values of $k$, showing comparable performances with respect to baselines.
predicted activity score average seed score
Mol-CycleGAN 0.381 0.179
Gargoyles 0.782 0.122
Pair-tuning 0.844 0.007
Table 8: Performance on the unconstrained DRD2 activity optimization task with respect to the initial value. Baseline performances are reported from Erikawa et al. (2023)[34].
DRD2 activity optimization
DRD2 activity optimization results are reported in Table 8. Activity scores are calculated using the trained predictor from Olivecrona et al. (2017)[35]. Average activity scores of the input seed molecules are also shown. Different baseline performances are reported using different test seed molecules, and we are interested in comparing the activity improvement of an experiment with the set of test seed molecules. For pair-tuning, performance reported considers $k=20$ generations per seed molecule and we use the top 1 for each, similar to the baselines, Mol-CycleGAN [36] — a graph-based generation method that uses a CycleGAN optimization scheme — and Gargoyles [34], which uses Monte Carlo Tree Search in fragment space to optimize molecular graphs based on an evaluation function. Results show that pair-tuning generates molecules with the highest activity improvement with respect to the seed molecules.
Conclusion
In this work, we investigate the effect of training data bias and diversity on the downstream performances of a generative chemical language model named GP-MoLFormer, built on top of a recent transformer architecture for chemical language modeling. We show how bias in training data can induce memorization and thus impact the novelty of generation. To our knowledge, this is the first report demonstrating training data memorization and its impact on performance of a generative chemical language model pre-trained on billion-scale data. We further analyze how the commonly used metrics for comparing models’ generations with a held-out test set are affected by the diversity of the training distribution; hence, the sizes of the test set and the generation set should be carefully considered before making a conclusion based on those metrics. Finally, we show the generality of the proposed GP-MoLFormer architecture on two distinct targeted design tasks, i.e., scaffold-constrained molecular decoration and unconstrained property-guided molecular optimization, where the model performs better or on par when compared to baselines.
Methods
Model Details
The GP-MoLFormer decoder uses the transformer block used in MoLFormer. To avoid the quadratic complexity associated with regular attention computations in vanilla transformers [37], MoLFormer utilized a base 12-layer transformer architecture with linear attention [12], wherein each layer has $12$ attention heads and a hidden state size of $768$. A Generalized Feature map [12] for the linear attention was chosen.
To better model positional dependence of tokens within a SMILES string, MoLFormer deviates from using the default absolute position embeddings and instead uses rotary embeddings [13]:
$\textnormal{Attention}_{m}(Q,K,V)=\frac{\sum_{n=1}^{N}\langle\varphi(R_{m}q_{m}),\varphi(R_{n}k_{n})\rangle v_{n}}{\sum_{n=1}^{N}\langle\varphi(R_{m}q_{m}),\varphi(R_{n}k_{n})\rangle},$
where $Q,K,V$ are the query, key, and value respectively, and $\varphi$ a random feature map. The GP-MoLFormer is trained on the next token prediction task using a cross-entropy objective: $L_{LM}(w1,...,wn)=\sum_{i}{\log{P}(w_{i}|w_{j<i})}$.
Datasets and Tokenization
We used two datasets for pre-training by combining SMILES from the PubChem [38] and ZINC [9] databases with varying proportions from each. The Raw dataset contains a total of 1.1B SMILES strings;  [^1] M of them are from the PubChem dataset, whereas the larger 1B portion comes from the ZINC database. The Clean dataset is a de-duplicated version of the Raw dataset, which comprises 650M SMILES. We utilized the tokenizer from Schwaller et al. (2019)[39] to construct a vocabulary. All SMILES sequences from both PubChem and ZINC are converted to a canonical format with no isomeric information using RDKit [40] followed by de-duplication (Clean only) and tokenization. All unique tokens extracted from the resulting output give us a vocabulary of $2357$ tokens plus $5$ special tokens, resulting in a total of $2362$ vocabulary tokens which are used for all pre-trained models considered in this paper, irrespective of pre-training dataset size. The post tokenization sequence length of the molecules range from $1$ to just over $2000$ tokens. We decide to restrict the sequence length range from $1$ token to $202$ tokens, special tokens inclusive, to reduce computation time. Since over $99.4$ percent of all molecules from our dataset contain less than $202$ tokens, we hypothesize that the removal of molecules with more than $202$ tokens would be of minimal negative impact on pre-training.
Large Scale Training and Parallelization
For pre-training, we use the causal language model objective defined in Devlin et al. (2019)[41]. The training was performed for 4 epochs through the entire combined PubChem+ZINC dataset with a fixed learning rate of $1.6\times 10^{-4}$ and a batch size of $1600$ molecules per GPU on a total of $16$ A100 80 GB GPUs over $2$ servers connected via EDR Infiniband fabric. It should be noted that as the number of GPUs utilized increased, we found an increase in learning rate was necessary up to a factor of $8$.
In order to scale our training to large datasets (>1B data points), we relied on adaptive bucketing of mini-batches by sequence length, as well as parallelization via distributed data-parallel training. The combination of linear attention, bucketing, and data parallelism allowed us to reduce the number of GPUs needed from roughly 1000 for quadratic attention with no bucketing to 16.
Pair-tuning
The paired molecule datasets for pair-tuning experiments were taken from Jin et al. (2019)[42]; The QED paired data used for training consists of 70644 molecule pairs where the first/seed molecule has a QED value in the range of 0.7–0.8 while the second/target molecule has a QED of 0.9–1.0. The penalized logP paired data consists of 60227 molecule pairs. It should be noted that while the paired datasets were collected such that molecular similarity within the pair is 0.4 and 0.6 for QED and logP, respectively, we demonstrate pair-tuning only on unconstrained property optimization tasks — we do not account for similarity preservation. The test set size for both QED and penalized logP optimization was 800. For the DRD2 binding optimization task, we used 34,404 molecule pairs from ZINC and Olivecrona et al. (2017)[35] for training and a test set of 1000 molecules[42]. For scoring the generated molecules, the bioactivity prediction model from Olivecrona et al. (2017)[35] is used; inactive compounds were defined with $p<0.05$ and actives were with $p>0.5$.
The vocabulary includes 20 randomly initialized prompt embeddings as well as the <unk> embedding from GP-MoLFormer training. For training, we prepended all 20 prompt embeddings to the <bos> embedding, followed by the embeddings of the first/seed molecule in a specific pair. We then add the <unk> embedding at the end of the first/seed molecule. After the <unk> embedding, we add the embeddings of the target molecule, followed by the <eos> embedding.
For evaluation, we do a forward pass using the following sequence: the first 20 prompt embeddings + the <bos> embedding + the input molecule embeddings + the <unk> embedding. After that, we sample from the token distribution generated by GP-MoLFormer until <eos> is encountered. For all pair-tuning experiments, batch size was set to 35, the learning rate was fixed at $3\times 10^{-2}$, and the number of epochs run was 1000. Each epoch took 6 minutes to complete on a single GPU. During pairtuning generation, GP-MoLFormer takes $0.00375$ seconds for a single forward pass.
Supplementary Information
Related Work
Earlier work on de novo molecule generation mapped a SMILES representation of a molecule into a continuous latent space using character based recurrent neural networks [14] and variational autoencoders (VAE)[43]. The VAE models have an encoder that converts the discrete molecular representation of a molecule into a continuous representation and a decoder that converts the continuous representation back to a discrete representation. The continuous representation allows for efficient de novo generation of molecules by sampling from the latent space. A property predictor is often trained on the latent embeddings from the encoder and is used to search for compounds that have specific desirable properties (conditional generation) using gradient-based or Bayesian optimization techniques.
However, there are several drawbacks for this approach. One drawback is that the decoded SMILES string may not always represent a valid molecule, often violating chemical constraints like valency. To alleviate this problem, several approaches have been proposed to augment the model with rules or mechanisms to ensure that the decoded SMILES strings are always valid [19, 44, 45, 46]. Another solution is to use alternate more robust string representations like DeepSMILES [47] SELFIES[23] and Group SELFIES [48]. While SELFIES-based representation ensures that all the decoded molecules are structurally valid, there is recent work that indicates that invalid SMILES are low-likelihood samples and are beneficial for chemical language models to better model the input molecule distribution [24].
Another drawback of this approach is that the property predictors trained on the latent embeddings are not as accurate as predictors trained end-to-end on the molecule itself. This necessitates a post-processing step, where the generated molecules have to be filtered using another, more accurate, predictor [49]. LIMO [20] avoids this problem by freezing the encoder and using a property predictor trained directly on the discrete representation of a molecule to optimize the latent space using an inceptionism-like approach.
Along with string generation methods, several graph and topological generative models have also been proposed [33, 50, 51, 52, 53, 32, 31, 54]. Another approach to molecule optimization is to use reinforcement learning to generate or modify a molecular graph [31, 30, 55, 35, 33]. While these methods have been shown to be successful against a variety of molecular optimization tasks, they often require multiple calls to the property predictor during the optimization process and hence are difficult to use along with computationally expensive property predictors.